Project

General

Profile

11302013

Nightly Observing Plan for MJD 56627 (starting evening of Nov 30, 2013)

Observers

Observer 1 Observer 2 Run Manager SISPI Expert on Call
Rutu Das Lyndsay Old Kevin Reil
ext. 521 ext. 516 ext. 513 or 630 338 9267

CTIO staff (Telescope Operators)

Day Support Night Assistant
Hernan Tirado, Javier Rojas, Humbert Orrego, Marco Bonati Claudio Aguilera

DECam Configuration and Setup

SISPI Configuration File DES.ini
DECam constants DECAM:CURRENT
Proposal ID 2012B-0001
Image Processing (Calibration) NCSA
Image Processing (First Cut) NCSA

Almanac
mjd 56627, date 20131130, Site CTIO, UT-offset 3 hr
Current local time 15:33 CLST     UT 18:33     SidTim 18:30
Sunset     20:38 CLST  23:38 UT   SidTim  354 deg  
Eve. -10   21:18 CLST  00:18 UT   SidTim    4 deg  Zeropoint; standards
Eve. -14   21:40 CLST  00:40 UT   SidTim    9 deg  Begin obstac
Eve. -18   22:04 CLST  01:04 UT   SidTim   15 deg  Last chance standards
Midpoint   01:32 CLST  04:32 UT   SidTim   67 deg  
Morn -18   05:00 CLST  08:00 UT   SidTim  119 deg  Begin astronomical twilight
Moonrise   05:22 CLST  08:22 UT   SidTim  125 deg  
Morn Std   05:23 CLST  08:23 UT   SidTim  125 deg  End obstac; begin standards
Morn -10   05:46 CLST  08:46 UT   SidTim  131 deg  
Sunrise    06:26 CLST  09:26 UT   SidTim  141 deg  
Moon position (deg): RA 222.4, Dec -14.9 Phase: 0.158

Weather Conditions

*Tololo Weather Links .

http://www.meteochile.gob.cl/cna/irsat_alta.png

Clear at 03:30 CLST

Observing Plan and Goals for the Night

  1. Afternoon calibrations script to use: prenight_calibration.json
  2. Start of night standards (preceed with zeros script).
  3. Nightly out of focus image. script to use: defocus_nightly.json (right after last standard, no break)
  4. ObsTac.
  5. End of night standards.
  6. End of night calibrations. if start of night calibrations good use: postnight_calibration.json, otherwise use postnight_zeros_flats.json

Start of Night Standards MANUAL SLEW TO THESE FIELDS Wind screen down.
UT Date = '2013/12/01 00:18' Local time = 21:18
SDSSJ0320-0000 ra 03:20:00 dec +00:00:00 secz 1.7 script to use: DES_std_sdssj0320_Yzirg
SDSSJ2140-0000 ra 21:40:00 dec +00:00:00 secz 1.5 script to use: DES_std_sdssj2140_Yzirg
SDSSJ0000-0000 ra 00:00:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj0000_Yzirg
nightly_defocus.json

End of Night Standards MANUAL SLEW TO THESE FIELDS Check Wind screen down.
UT Date = '2013/12/01 08:23' Local time = 05:23
SDSSJ1048-0000 ra 10:48:00 dec +00:00:00 secz 1.4 script to use: DES_std_sdssj1048_grizY
SDSSJ0843-0000 ra 08:43:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj0843_grizY

Notes from 4pm Meeting

Agenda: Lost clocks last night. Also lost S30. Data taking called off by run manager. Loose cables found today. As of 3:30 all sensors and backplanes look fine in bias images (except S30/N30).

Run Manager's Meeting 11/30/2013
---------------------------------
Alistair, Kevin, Klaus, Tom, Marco, Eric

Telescope
---------
Moving around for CCD tests.
Camera
------
Kevin:
Night before last, backplane 1 had a lot of noise. This had happened
before several times. Sol'n had been to Power Cycle (PC) the crate but that no longer
worked. Decided to take data with noisy (65 cts RMS) on bkpln 1. This really
is some nasty high frequency noise. Images after 258504 onwards
is OK and maybe a few before that.
Yesterday best guess was the Power Supply. Changed it. That didn't help.
JE -> MB suggested switching backplane that is master (origin of clocks).
No time for calib or standards.
1st image didn't deliver. Many backplanes missing. None of the
non master backplanes had clocks. (BKPLN 1 was master). PC didn't
help. Switched to BKPLN 3 as master - OK except noise on BKPLN 1
but S30 showed up dead on one side. S30 is on backplane 3.
Shift was sent to bed.
Marco:
Javier looked at synch cables and found 1 loose (1 --> 3). It must
have come completely loose yesterday afternoon. Several images w/
bckpln 1 as master showed no noise. Even with backplane 3 as master.
Now they are moving the telescope around to see if the noise comes
back. It could be the loose synch cable was the root of the problems.
Right now backplane 3 is the master.
S30 is a different story. They are still looking at it. One of the
clocks is at the wrong value (-4.5V instead of -12V). Voltage was fine at
6 PM yesterday but NG after dinner. There are some ideas for what might be
wrong. The tests involve testing the particular video board on the
bench. This is to be planned for next week. Also put off checking
if the particular bias (Vreset) behaves resistively. Juan and
Marco are making a careful plan.
Marco says Juan E. notes telemetry shows a change in electronics
Temp at the time of the power cycles.
If the noise comes back plan is to goto backplane 1 as master.
DESDM
-----
DQ
--
We are waiting to hear about the DQ on the recent data. This
maybe more complicated than the standard DQ tests can handle.
Might be we are into partial focal plane images. I don't know if
DESDM can handle that for a subset of images during a night.
Plan for Tonight
----------------
Weather looks fantastic. Likely to be SN deadman observing.
PTC tomorrow PROPID 2013a-9999.
AOB
---
Need a report on system 1.
Hexapod is working. Arm #6 is being controlled on the motor encoder.
ADS guys think it's not the incremental encoder that is bad.
Problem is likely to be one of the connections someplace. At some
suitable time there'll be some more tests.
We are now using the spare encoder distribution box, so that's not the
problem.
Kevin found 8 RTDs on the camera ready to deploy on the cage.
Mike Warner is getting ready to deploy the accelerometer in a few weeks.
3D anemometer is in transit to FNAL from OSU and it won't slow down
much.